OpenStudy (anonymous):

Match each equation with its equivalent equation. Match Term Definition four x equals sixteen A) two x equals eight x divided by twelve equals three B) x divided by four equals nine x divided by eight equals four C) x divided by four equals eight four x equals eight D) two x equals four

OpenStudy (anonymous):

pleaseee help!

OpenStudy (anonymous):

@texaschic101 !!!!

OpenStudy (anonymous):

@iceicebaby

OpenStudy (texaschic101):

4x = 16 --- reduce by dividing by 2 2x = 8 (A) x/12 = 3 ---x/4 = 9......the answers to both are : x = 36 (B) x/8 = 4 --- x/4 = 8...the same answer to both is : x = 32 (C) 4x = 8 --- reduce 2x = 4 (D)

OpenStudy (anonymous):

can you please help me with oje last one i cant get it @texaschic101